I am having trouble with this equation: 3x+4=2x-3
I don't know the formula. Where do I start to answer this question?
Thank you for spending your time to help me!

Answer by checkley71(8403)

3X+4=2x-3 or 3x-2x=-3-4 or x=-7 proof
3(-7)+4=2(-7)-3 or -21+4=-14-3 or -17=-17